# SSD4Rec: A Structured State Space Duality Model for Efficient Sequential Recommendation

### v1

@@ 2.4.3. Prediction Layer
Finally, in the prediction layer, SSD4Rec using the predicted representations to generate the final output prediction scores. In a specific batch $b$ with $K$ users, the prediction process can be written by
(12) $\hat{\mathcal{Y}}_{b}=\text{Softmax}(\mathcal{P}_{b}*\mathcal{Q}),$
@@ 2.5. Learning Objective
To learn the parameters of the proposed method SSD4Rec for sequential recommendation, we use a representative objective function, Cross-Entropy (CE) loss, between the predicted probability distributions and the classification labels of corresponding positive items for optimizing the model. Mathmatically, in a batch with $K$ users, the loss function can be expressed as
@@
2.6. Time Complexity Analysis
@@
SASRec (Kang and McAuley, 2018) Mamba4Rec (Liu et al., 2024) SSD4Rec
@@ 3.1.1. Datasets
To showcase the efficiency and effectiveness of our proposed approach, we perform extensive experiments across four benchmark datasets: Amazon-Beauty (referred to as Beauty), Amazon-Video-Games (referred to as Games), LastFM, and MovieLens 1M (abbreviated as ML1M). The initial two datasets, sourced from the Amazon e-commerce platform [^1] https://nijianmo.github.io/amazon/, encapsulate a wide array of user engagements with Beauty and Video Games products. Moreover, the ML1M [^2] https://grouplens.org/datasets/movielens/1m/ dataset comprises a compilation of movie ratings provided by users on the MovieLens platform. Finally, the KuaiRand dataset [^3] https://kuairand.com/, retrieved from the recommendation logs of the video-sharing mobile application Kuaishou, includes millions of interactions involving randomly exposed items. The statistics for these four datasets are outlined in Table 2. We follow the leave-one-out policy (Kang and McAuley, 2018) for training-validation-testing partition.
@@ 3.1.2. Baselines
To demonstrate the superiority of the proposed method, four representative sequential recommendation methods and a recently developed Mamba-based method are used as baselines, namely Caser (Tang and Wang, 2018): a typical convolution-based sequential recommender; GRU4Rec (Hidasi et al., 2015): a RNN-based method constructed by Gated Recurrent Units; BERT4Rec (Sun et al., 2019): a sequential recommender system based on bidirectional Transformers trained using the BERT-style cloze task; SASRec (Kang and McAuley, 2018): a representative approach for sequential recommendation leveraging self-attention mechanisms; and Mamba4Rec (Liu et al., 2024): a pioneering model constructed by SSMs.
@@ 3.1.4. Hyper-parameter Settings
Our evaluation is implemented based on PyTorch  [^4] https://pytorch.org/ and RecBole  [^5] https://github.com/RUCAIBox/RecBole. First, to initialize the proposed Mamba-based block, we designate the model dimension as 256 and the SSM state expansion factor as 64. Second, the weighted indicator for backward SSD computation $\beta$ is searched from 0.1 to 0.9 in 0.2 increments. Moreover, the search space for the learning rate includes values of {0.0001, 0.001, 0.01, 0.1}, while for the batch size, we consider {64, 256, 512, 1024, 2048}. Fourth, the number of Bi-SSD cells is searched from 1 to 3. To accommodate the baselines, we set fixed sequence lengths of 200 for ML1M and 50 for other datasets. Furthermore, for a fair comparison, we ensure consistency in the key hyperparameters across models while setting the other default hyperparameters for baseline methods as suggested in their respective papers. Finally, we employ the Adam (Kingma, 2014) optimizer in a mini-batch manner.
@@ 3.2. Recommendation Performance Comparison
We first compare the recommendation performance between the proposed model and all baselines over the four benchmark datasets. Table 3 and Table 4 showcase the overall evaluation results. From the two tables, we can make the following observations.
  • Our proposed SSD4Rec consistently outperforms all the baselines across all datasets regarding NDCG@10&20 and MRR@10&20. On average, SSD4Rec significantly exceeds the strongest baselines by 6.17% on NDCG@10 and 7.73% on MRR@10 in the KuaiRand dataset. Such improvement demonstrates the effectiveness of our proposed method and the great potential of exploring the Structured State Space Duality property in sequential recommendations.
@@
Table 3. Sequential recommendation performance on the ML1M and Beauty datasets, where NG@K refs to NDCG@K, and the best and second-best results are bold and underlined, respectively.
Method ML1M Beauty
NG@10 NG@20 MRR@10 MRR@20 HR@10 HR@20 NG@10 NG@20 MRR@10 MRR@20 HR@10 HR@20
Caser 0.1714 0.1948 0.1354 0.1418 0.2892 0.382 0.0294 0.0355 0.0222 0.0239 0.0531 0.0774
GRU4Rec 0.1655 0.1906 0.1276 0.1344 0.2901 0.3897 0.0307 0.0386 0.0221 0.0242 0.0593 0.0907
BERT4Rec 0.1846 0.2082 0.1441 0.1506 0.3171 0.4103 0.0219 0.0279 0.0159 0.0175 0.0419 0.0656
SASRec 0.1718 0.1963 0.1320 0.1386 0.3023 0.4000 0.0395 0.0480 0.0261 0.0284 0.0829 0.1168
Mamba4Rec 0.1847 0.2094 0.1456 0.1523 0.3124 0.4103 0.0442 0.0501 0.0361 0.0376 0.0709 0.0942
SSD4Rec 0.1889 0.2145 0.1495 0.1565 0.3152 0.4194 0.0452 0.0520 0.0363 0.0382 0.0745 0.1015
Table 4. Sequential recommendation performance on the Games and KuaiRand datasets, where NG@K refs to NDCG@K, and the best and second-best results are bold and underlined, respectively.
Method Games KuaiRand
NG@10 NG@20 MRR@10 MRR@20 HR@10 HR@20 NG@10 NG@20 MRR@10 MRR@20 HR@10 HR@20
Caser 0.0460 0.0587 0.0330 0.0364 0.0891 0.1396 0.0545 0.0692 0.0414 0.0453 0.0982 0.1571
GRU4Rec 0.0520 0.0654 0.0373 0.0409 0.1011 0.1543 0.0563 0.0722 0.0426 0.0470 0.1017 0.1653
BERT4Rec 0.0372 0.0482 0.0263 0.0293 0.0735 0.1170 0.0534 0.0683 0.0404 0.0444 0.0968 0.1563
SASRec 0.0531 0.0689 0.0330 0.0373 0.1189 0.1817 0.0567 0.0733 0.0426 0.0471 0.1040 0.1705
Mamba4Rec 0.0605 0.0738 0.0461 0.0497 0.1081 0.1610 0.0558 0.0710 0.0427 0.0468 0.0994 0.1601
SSD4Rec 0.0610 0.0740 0.0461 0.0497 0.1102 0.1617 0.0602 0.0759 0.0460 0.0503 0.1076 0.1704
Table 5. Comparison of the computational efficiency on the ML1M dataset, which involves measuring the training time per epoch and the inference time for all test users in seconds (s).
L=200 N SASRec* Mamba4Rec SSD4Rec Imp.* N=256 L SASRec* Mamba4Rec SSD4Rec Imp.*
Training (s) 32 106.18 44.98 22.90 363.67% Training (s) 50 42.79 119.87 29.37 45.69%
64 112.43 82.16 24.69 355.37% 100 69.50 173.78 47.73 45.61%
128 125.24 156.45 35.86 249.25% 200 161.91 326.87 74.13 118.41%
256 161.91 326.87 74.13 118.41% 400 450.64 687.46 101.51 343.94%
Inference (s) 32 0.158 0.061 0.039 305.13% Inference (s) 50 0.057 0.217 0.057 0.00%
64 0.178 0.114 0.046 286.96% 100 0.111 0.272 0.089 24.72%
128 0.199 0.218 0.062 220.97% 200 0.27 0.484 0.129 109.30%
256 0.27 0.484 0.129 109.30% 400 0.75 1.152 0.167 349.10%
@@ 3.3. Efficiency Analysis
In this subsection, we analyze the training and inference efficiency of SSD4Rec compared with two representative sequential recommendation methods, i.e., SASRec and Mamba4Rec. The evaluation results on the ML1M dataset are presented in Table 5. We can see that the two Mamba-based models, i.e., Mamba4Rec and SSD4Rec are much more efficient compared to the Attention-based SASRec model when $N<L=200$, which in line with the discussion in Section 2.6, i.e., Mamba enjoys linear scaling in sequence length. When presented with inputs of substantial dimensions, such as $N=256$, Mamba4Rec exhibits lower efficiency compared to SASRec, even when $L=400>N$, showcasing the constraints of Mamba-1 in managing high-dimensional inputs (Dao and Gu, 2024). In contract, SSD4Rec achieves superior computational efficiency to SASRec and Mamba4Rec in all the cases. This superior performance can be attributed to the utilization of the novel SSD layer and the advancing hardware-aware computation algorithm (i.e., the block-based matrix multiplication). To be specific, when given a lengthy interaction list, i.e., $N=32$ and $L=200$, the proposed model is 3.6$\times$ and 3$\times$ faster than the Attention-based method SASRec in the training and inference phases, respectively. This substantial acceleration underscores a notable enhancement in computational efficiency. Moreover, even when configured with high-dimensional input sequences, i.e., $N=256>L=200$, our model experiences a speedup of exceeding 100% in comparison to SASRec for both the training and inference tasks, underscoring the significance of the hardware-aware computation algorithm employed by SSD.
Table 6. Ablation studies conducted by eliminating the effects of each key component, where NG@10 refs to NDCG@10.
@@
NG@10 MRR@10 NG@10 MRR@10
Full 0.1889 0.1495 0.0602 0.0460
w/o Bidirection 0.1869 0.1474 0.0593 0.0454
w/o Register 0.1871 0.1466 0.0545 0.0421
w/o Varlen 0.1852 0.1456 0.0565 0.0434
w PE 0.1736 0.1353 0.0566 0.0431
1 Layer 0.1872 0.1483 0.0579 0.0441
3 Layers 0.1933 0.1529 0.0597 0.0455

### v2

@@ 2.4.3. Prediction Layer
Given the integrated input $\mathcal{X}=\{\mathcal{E}_{b},\mathcal{R}_{b}\}$ associated with $K$ users in the $b$-th batch, after $M$ Bi-SSD layers, we get the hidden representations of user-item sequence $\mathcal{H}^{M}$. Notably, in the final Bi-SSD layer, we utilize the generated representation of the last item for each user to serve as the predicted preference representation. This process can be formulated as follows:
(22) $\mathcal{H}^{M}=[\mathbf{H}_{u_{1}},...,\mathbf{H}_{u_{k}},...,\mathbf{H}_{u_{K}}]\rightarrow\mathcal{P}_{b}=\{\hat{\mathbf{p}}_{1},...,\hat{\mathbf{p}}_{k},...,\hat{\mathbf{p}}_{K}\},$
where $\mathbf{H}_{u_{k}}$ denotes the sequence of hidden representations of user $u_{k}$ in $\mathcal{H}^{M}$. $\hat{\mathbf{p}}_{k}\in\mathbb{R}^{D}$ corresponds to the representation of the last item associated with user $u_{k}$ in $\mathbf{H}_{u_{k}}$ and acts as the predicted representation. Using these predicted representations in a batch $\mathcal{P}_{b}$, SSD4Rec can calculate the final output prediction scores. In a specific batch $b$ with $K$ users, the prediction process can be written by
(23) $\hat{\mathcal{Y}}_{b}=\text{Softmax}(\mathcal{P}_{b}*\mathcal{Q}).$
@@ 2.5. Learning Objective
To learn the parameters of the proposed method SSD4Rec for the sequential recommendation, we use a representative objective function, Cross-Entropy (CE) loss, between the predicted probability distributions and the classification labels of corresponding positive items for optimizing the model. Mathematically, in a batch with $K$ users, the loss function can be expressed as
@@
2.6. Discussion
2.6.1. Bidirectional Learning Analysis
It is not straightforward and intuitive for conventional architectures to train the bidirectional model for a sequential recommendation. For instance, jointly conditioning on both left and right context in a deep bidirectional Transformer-based model would cause information leakage, i.e., allowing each item to indirectly “see the target item”. This could make predicting the future trivial, and the network would not learn anything useful (Sun et al., 2019). To avoid information leakage, the proposed SSD4Rec conducts the reverse process solely on the previous items based on the property of Sequence Alignment (Hwang et al., 2024) in the SSD architecture. Together with the variable-length input, our approach implements data augmentation by creating separate samples. This means for each individual sample with a target item $v_{L}$, we have the forward item sequence $\{v_{1},v_{2},...,v_{L-1}\}$ and the backward item sequence $\{v_{L-1},...,v_{2},v_{1}\}$ as model inputs. Although it may result in a slight decrease in training efficiency, one notable advantage of this approach is its ability to enhance the diversity of samples in each batch, unlike the Transformer-based learning process (Kang and McAuley, 2018), where data-augmented samples of the same user are bound to be present in the same batch.
2.6.2. Time Complexity Analysis
@@
SASRec (Kang and McAuley, 2018) Mamba4Rec (Liu et al., 2024a) SSD4Rec
@@ 3.1.1. Datasets
To showcase the efficiency and effectiveness of our proposed approach, we perform extensive experiments across four benchmark datasets: Amazon-Beauty (referred to as Beauty), Amazon-Video-Games (referred to as Games), MovieLens 1M (abbreviated as ML1M), and KuaiRand. The initial two datasets, sourced from the Amazon e-commerce platform [^1] https://nijianmo.github.io/amazon/, encapsulate a wide array of user engagements with Beauty and Video Games products. Moreover, the ML1M [^2] https://grouplens.org/datasets/movielens/1m/ dataset comprises a compilation of movie ratings provided by users on the MovieLens platform. Finally, the KuaiRand dataset [^3] https://kuairand.com/, retrieved from the recommendation logs of the video-sharing mobile application Kuaishou, includes millions of interactions involving randomly exposed items. Notably, we filter out users and items with fewer than five interactions in each dataset. The statistics for these four datasets are outlined in Table 2. We follow the leave-one-out policy (Kang and McAuley, 2018) for training-validation-testing partition.
@@ 3.1.2. Baselines
To demonstrate the superiority of the proposed method, four representative sequential recommendation methods and two recently developed Mamba-based methods are used as baselines, namely:
  • Caser (Tang and Wang, 2018): a typical convolution-based sequential recommender that integrates Convolutional Neural Networks (CNNs) to capture sequential features in historical user-item interactions;
  • GRU4Rec (Hidasi et al., 2015): a classical recurrent-based method in sequential recommendation, which is constructed by Recurrent Neural Networks (RNNs) with Gated Recurrent Units;
  • BERT4Rec (Sun et al., 2019): a sequential recommender system based on bidirectional Transformer layers trained using the BERT-style cloze task;
  • SASRec (Kang and McAuley, 2018): a representative approach for sequential recommendations that utilize self-attention mechanisms to enhance its modeling capabilities;
  • Mamba4Rec (Liu et al., 2024a): a pioneering model built upon the novel Structured State Space Model, namely Mamba;
  • SIGMA (Liu et al., 2024b): a hybrid sequential recommendation model based on Mamba and GRU, which enhances context modeling capability by introducing Partially Flipped Mamba with a Dense Selective Gate.
@@ 3.1.4. Hyper-parameter Settings
Our evaluation is implemented based on PyTorch  [^4] https://pytorch.org/ and RecBole  [^5] https://github.com/RUCAIBox/RecBole. First, to initialize the proposed Mamba-based block, we designate the model dimension as 256 and the SSM state expansion factor as 64. Second, we employ Sequential Grid Search to sequentially determine the two critical hyper-parameters for our model: the backward indicator $\beta$ in Eq. (14) and the mask ratio $\rho$ in Eq. (3). The search ranges for both parameters are defined as {0, 0.1, 0.2, 0.3, 0.5, 0.7, 0.9}. Third, the search space for the learning rate includes values of {0.0001, 0.001, 0.01, 0.1}, while for the batch size, we consider {64, 256, 512, 1024, 2048}. The final values for these two hyper-parameters (learning rate and batch size) are set to 0.001 and 1024, respectively. Fourth, the number of Bi-SSD cells is explored within the range of 1 to 3. To accommodate the baselines, we set fixed sequence lengths of 200 for ML1M and 50 for other datasets. Due to the sparsity of the Amazon-beauty and Amazon-games datasets, their dropout rate is set to 0.4. In contrast, the dropout rates for MovieLens-1M and KuaiRand are set at 0.2. Furthermore, for a fair comparison, we ensure consistency in the key hyperparameters across models while setting the other default hyperparameters for baseline methods as suggested in their respective papers. Finally, all experiments are conducted on a single NVIDIA 4090 GPU, and we employ the Adam (Kingma, 2014) optimizer in a mini-batch manner. The running configuration and settings are listed in Table 3.
Table 3. Hyperparameter Settings corresponding to different datasets
Datasets $\beta_{best}$ $\rho_{best}$ $L$ Dropout # Bi-SSD Layers Batch Size Learning Rate
ML1M 0.1 0.1 200 0.2 2 1024 0.001
Beauty 0.1 0.2 50 0.4 1 1024 0.001
Games 0.1 0.1 50 0.4 1 1024 0.001
KuaiRand 0.1 0.2 50 0.2 2 1024 0.001
Range [0,1) [0,1) - - {1,2,3} {4, 256, 512, 1024, 2048} {0.0001, 0.001, 0.01, 0.1}
@@ 3.2. Recommendation Performance Comparison
We first compare the recommendation performance between the proposed model and all baselines over the four benchmark datasets. Table 4 showcases the overall evaluation results. From the table, we can make the following observations.
  • Our proposed SSD4Rec consistently outperforms all the baselines across all datasets regarding NDCG@10&20 and MRR@10&20. On average, SSD4Rec significantly exceeds the strongest baselines by 13.80% on NDCG@10 and 16.14% on MRR@10 in the ML1M dataset. Such improvement demonstrates the effectiveness of our proposed method and the great potential of exploring the Structured State Space Duality property in sequential recommendations.
@@
Table 4. Sequential recommendation performance on four different datasets, where Imp.* denotes the relative performance improvement in comparison to the specified baseline marked by *. The best and second-best results are bold and underlined, respectively.
Dataset Metric Caser GRU4Rec BERT4Rec SASRec* Mamba4Rec SIGMA SSD4Rec Imp.*
ML1M NDCG@10 0.1714 0.1655 0.1846 0.1718 0.1847 0.1777 0.1955 13.80%
NDCG@20 0.1948 0.1906 0.2082 0.1963 0.2094 0.2026 0.2225 13.35%
MRR@10 0.1354 0.1276 0.1441 0.1320 0.1456 0.1361 0.1533 16.14%
MRR@20 0.1418 0.1344 0.1506 0.1386 0.1523 0.1431 0.1607 15.95%
HR@10 0.2892 0.2901 0.3171 0.3023 0.3124 0.3136 0.3334 10.29%
HR@20 0.3820 0.3897 0.4103 0.4000 0.4103 0.4121 0.4406 10.15%
Beauty NDCG@10 0.0294 0.0307 0.0219 0.0395 0.0442 0.0376 0.0470 18.99%
NDCG@20 0.0355 0.0386 0.0279 0.0480 0.0501 0.0455 0.0535 11.46%
MRR@10 0.0222 0.0221 0.0159 0.0261 0.0361 0.0284 0.0383 46.74%
MRR@20 0.0239 0.0242 0.0175 0.0284 0.0376 0.0305 0.0400 40.85%
HR@10 0.0531 0.0593 0.0419 0.0829 0.0709 0.0678 0.0756 -8.81%
HR@20 0.0774 0.0907 0.0656 0.1168 0.0942 0.0994 0.1016 -13.01%
Games NDCG@10 0.0460 0.0520 0.0372 0.0531 0.0605 0.0592 0.0611 15.07%
NDCG@20 0.0587 0.0654 0.0482 0.0689 0.0738 0.0746 0.0747 8.41%
MRR@10 0.0330 0.0373 0.0263 0.0330 0.0461 0.043 0.0464 40.61%
MRR@20 0.0364 0.0409 0.0293 0.0373 0.0497 0.0472 0.0500 34.05%
HR@10 0.0891 0.1011 0.0735 0.1189 0.1081 0.1342 0.1099 -7.57%
HR@20 0.1396 0.1543 0.1170 0.1817 0.1610 0.2002 0.1635 -10.02%
KuaiRand NDCG@10 0.0545 0.0563 0.0534 0.0567 0.0558 0.0572 0.0593 4.59%
NDCG@20 0.0692 0.0722 0.0683 0.0733 0.0710 0.0727 0.0757 3.27%
MRR@10 0.0414 0.0426 0.0404 0.0426 0.0427 0.0433 0.0448 5.16%
MRR@20 0.0453 0.0470 0.0444 0.0471 0.0468 0.0475 0.0493 4.67%
HR@10 0.0982 0.1017 0.0968 0.1040 0.0994 0.1036 0.1075 3.37%
HR@20 0.1571 0.1653 0.1563 0.1705 0.1601 0.1655 0.1731 1.52%
Table 5. Comparison of the computational efficiency on the ML1M dataset, which involves measuring the training time per epoch for all test users in seconds (s), where Imp.* denotes the relative computational speed improvement in comparison to the specified baseline marked by *; and SSD4Rec- represents the proposed model without the backward SSD block. The results indicate that SSD4Rec accelerates the training efficiency of the typical Attention-based and Mamba-based methods (i.e., SASRec and Mamba4Rec) by an average of 84.18% and 88.53%, respectively.
ML1M Training time
FIX $N/L$ SASRec* Mamba4Rec SSD4Rec- SSD4Rec Imp.*
$L=200$ 32 88.16s 32.92s 25.85s 43.06s 104.71%
64 91.93s 52.06s 26.91s 43.86s 109.57%
128 102.48s 90.76s 27.92s 49.13s 108.58%
256 134.29s 172.71s 45.56s 79.71s 68.49%
$N=256$ 50 38.53s 51.17s 23.79s 39.30s -1.96%
100 53.08s 66.59s 28.42s 47.29s 12.24%
200 134.29s 172.71s 45.56s 79.71s 68.49%
400 349.94s 489.20s 87.60s 115.38s 203.27%
Table 6. Comparison of the computational efficiency on the ML1M dataset, which involves measuring the inference time for all test users in millisecond (ms), where Imp.* denotes the relative computational speed improvement in comparison to the specified baseline marked by *; and SSD4Rec- represents the proposed model without the backward SSD block. The results indicate that SSD4Rec surpasses the inference efficiency of the typical Attention-based and Mamba-based methods (i.e., SASRec and Mamba4Rec) by an average of 72.35% and 53.63%, respectively.
ML1M Inference time
FIX $N/L$ SASRec* Mamba4Rec SSD4Rec- SSD4Rec Imp.*
$L=200$ 32 116ms 40ms 34ms 58ms 100.00%
64 126ms 58ms 38ms 53ms 137.74%
128 139ms 93ms 41ms 66ms 110.61%
256 176ms 180ms 68ms 122ms 44.26%
$N=256$ 50 33ms 82ms 29ms 47ms -29.79%
100 71ms 100ms 45ms 79ms -10.13%
200 176ms 180ms 68ms 122ms 44.26%
400 496ms 552ms 96ms 176ms 181.81%
@@ 3.3. Efficiency Analysis
In this subsection, we will mainly analyze the training and inference efficiency of SSD4Rec compared with two representative sequential recommendation methods, i.e., SASRec and Mamba4Rec. For fairness, it should be noted that the number of network layers for SSD4Rec is limited to 1. This corresponds with the single-layer structures of SASRec and Mamba4Rec. Due to limitations in GPU memory, we set both the training and evaluation batch sizes to 512 to prevent possible out-of-memory issues when N=256 and L=400. For a comprehensive evaluation across diverse sequence length $L$ and token dimension $N$, we fix one of the two parameters while varying the other. We also incorporate SSD4Rec- as a baseline in our efficiency analysis for training and inference time, which simply removes the backward SSD block structure from our processed SSD4Rec model. The evaluation results on the ML1M dataset are presented in Table 6. We can see that the two Mamba-based models, i.e., Mamba4Rec and SSD4Rec are much more efficient compared to the Attention-based SASRec model when $N<L=200$, which in line with the discussion in Section 2.6.2, i.e., Mamba enjoys linear scaling in sequence length. When presented with inputs of substantial dimensions, such as $N=256$, Mamba4Rec exhibits lower efficiency compared to SASRec, even when $L=400>N$, showcasing the constraints of Mamba-1 in managing high-dimensional inputs (Dao and Gu, 2024). In contrast, SSD4Rec achieves superior computational efficiency to SASRec and Mamba4Rec in most cases. Specifically, SSD4Rec only showcases slightly inferior training and inference speeds compared to the baselines when $N$ is significantly larger than $L$ (i.e., $N=256$ and $L=50/100$) However, it demonstrates a clear advantage as $L$ increases due to its linear computational complexity with respect to sequence length, as discussed in Table 1. Moreover, this superior performance can be attributed to the utilization of the advancing hardware-aware computation algorithm (i.e., the block-based matrix multiplication (Dao and Gu, 2024)) in our SSD-based model. For instance, when given a lengthy interaction list, i.e., $N=32$ and $L=200$, the proposed model is 104% and 100% faster than the Attention-based method SASRec in the training and inference phases, respectively. This substantial acceleration underscores a notable enhancement in computational efficiency. Even when configured with high-dimensional input sequences, i.e., $N=256>L=200$, our model experiences a speedup of exceeding 40% in comparison to SASRec for both the training and inference tasks. Moreover, the SSD4Rec- model without backward SSD block demonstrates superior computational efficiency compared to SASRec across all scenarios, including cases where $L$ is significantly smaller than $N$. These two observations highlight the substantial performance gains achieved through the advanced hardware-aware algorithm integrated within the SSD block.
Table 7. Ablation studies conducted by eliminating the effects of each key component.
@@
NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10 HR@10
Full 0.1955 0.1533 0.3334 0.0593 0.0448 0.1075
w/o Bidirection 0.1921 0.1521 0.3222 0.0585 0.0446 0.1049
w/o mask 0.1897 0.1488 0.3235 0.0581 0.0446 0.1034
w/o Varlen 0.1891 0.1484 0.3222 0.0572 0.0436 0.1025
w/o Register 0.1920 0.1509 0.3262 0.0565 0.0429 0.1022
w PE 0.1901 0.1496 0.3225 0.0585 0.0447 0.1049
1 Layer 0.1930 0.1523 0.3258 0.0584 0.0442 0.1055
3 Layers 0.1926 0.1505 0.3303 0.0557 0.0416 0.1020